# Existence of the Chapman-Enskog solution and its relation with first-order dissipative fluid theories

### v1

@@ I Introduction and preliminaries
In the relativistic scenario, the CE approximation was applied by pioneers in the field such as W. Israel Israel63 and S. R. de Groot deGrootLibro . However, the fact that its traditional form predicts first-order couplings between dissipative fluxes and spatial gradients of the state variables leads to its partial dismissal, based on the findings that such relations give rise to pathological theories for which the equilibrium configurations are generically unstable Hiscock . Recently, a new family of first-order theories has been proposed and shown to contain promising candidates to describe high temperature gases in curved spacetimes. These new theories allow one to consider a general frame and representation in which constitutive equations feature all possible couplings, including relations which involve time derivatives of the state variables. Under suitable restrictions, these theories have been shown to lead to physically sound equations BDNK_PRX ; pK19 . However, despite recent attempts bdnk19 ; rocha22 , their microscopical foundations are still not well understood from the point of view of the CE expansion. The purpose of this work is to shed new light on this problem.
@@
The second, more formal, method follows the lines of Ref. LaurePaper by projecting the source term $L_{F}\left[\ln f^{\left(0\right)}\right]$ onto the orthogonal complement of the kernel of $C$, which consists of the collision invariants. Since the operator $C$ is self-adjoint when defined in an appropriate Hilbert space, this assures that the source term lies in the image of $C$.
@@ II Non-relativistic case
In order to compare the order of magnitude of each term in Boltzmann’s equation (1) one introduces dimensionless variables, denoted by overbars, as follows: $t=\left(L/\left\langle c\right\rangle\right)\bar{t}$, $\vec{r}=L\vec{\bar{r}},\,$$\vec{v}=\left\langle c\right\rangle\vec{\bar{v}}$, $g=\sqrt{2}\left\langle c\right\rangle\bar{g}$, $f=\bar{f}/\left(\left\langle c\right\rangle^{3}L^{3}\right)$, $\sigma=\pi d^{2}\bar{\sigma}$ and $\vec{F}=\left(m\left\langle c\right\rangle^{2}/L\right)\vec{\bar{F}}$. Here $L$ is a characteristic macroscopic length scale, $\left\langle c\right\rangle$ a typical mesoscopic speed and $d$ a microscopic scale of the order of magnitude of the molecules’ diameter. Defining the mean free path as $\ell:=L^{3}/\left(\sqrt{2}\pi d^{2}\right)$ and introducing these scaled variables in Eq. (1) one obtains (omitting the bars)
@@ II.1 The traditional Chapman-Enskog procedure
where the quantities $\left(\mathcal{D}f\right)_{i}$ are still to be determined. This ansatz can be motivated by various arguments. The simplest one consists in noticing that the operator $\mathcal{D}$ involves only macroscopic variables, unlike the other two terms which depend on the molecular velocity. For a thorough discussion of this point the reader is referred to Refs. Ch-E and cercigniani1 for example.
@@
In order to determine $\left(\mathcal{D}n\right)_{0}$, $\left(\mathcal{D}\vec{u}\right)_{0}$ and $\left(\mathcal{D}T\right)_{0}$, we multiply both sides of Eq. (17) by the vector of collision invariants $\vec{\psi}=\left(1,\vec{v},\frac{1}{2}mv^{2}\right)^{T}$ and $f^{\left(0\right)}$ and integrate over momentum space, obtaining
@@ III Relativistic Case
The order of magnitude for the terms in Eq. (1) can be assessed in a similar way as in the non-relativistic scenario with the following additional assumptions: $p^{a}=m\left\langle c\right\rangle\bar{p}^{a}$, $f=\frac{1}{m^{3}\left\langle c\right\rangle^{3}L^{3}}\bar{f}$, $c=\langle c\rangle\bar{c}$, $q=e\bar{q}$, $F^{ab}=\frac{m\left\langle c\right\rangle^{2}}{Le}\bar{F}^{ab}$, $\mathcal{F}=\sqrt{2}m^{2}\left\langle c\right\rangle^{2}\bar{\mathcal{F}}$, where $e$ is the elementary charge. Proceeding as in Section II, one is led to the following dimensionless equation
@@
where we have once again omitted the bars over the dimensionless quantities in order to simplify the notation. Thus, in this scenario the collisional term also scales one order higher in $1/\text{Kn}$ than the transport part, which suggests the same expansion as in the previous case: $f=\sum_{i=0}^{\infty}\text{Kn}^{i}f^{\left(i\right)}$. As before, $f^{\left(0\right)}$ corresponds to a local equilibrium distribution function, parametrized by the dynamical variables $n$, $T$ and $u^{a}$ (where $u^{a}$ is the hydrodynamic four-velocity), and is given by the solution to $Q\left(f^{\left(0\right)}f^{\left(0\right)^{\prime}}\right)=0$ in Eq. (37), which leads to the Maxwell-Jüttner distribution function CercignaniKremer-Book :
@@
However, out of equilibrium, there is no unique way to associate the components of $N^{a}$ and $T^{ab}$ to the state variables and fluxes in the system. In fact, there is a large family of valid representations as explained in Refs. pK19 ; BDNK_PRX . In the present work the so-called Eckart (or particle) frame is considered. This particular choice is such that
@@ III.1 The relativistic generalization of the traditional Chapman-Enskog procedure
Notice that $u^{a}\frac{\partial f}{\partial x^{a}}$ in the first term on the right-hand side of Eq. (50) is the equivalent to the convective time derivative in the non-relativistic case. Thus it is reasonable to define
$\mathcal{D}f:=u^{a}\frac{\partial f}{\partial x^{a}},$ (51)
and to write the relativistic Liouville operator as
@@
In order to find expressions for $\left(\mathcal{D}n\right)_{0}$, $\left(\mathcal{D}T\right)_{0}$ and $\left(\mathcal{D}u^{c}\right)_{0}$ one can once again use the collision invariants. Indeed, multiplying Eq. (56) by $\vec{\psi}:=\left(1,p^{a}\right)^{T}$ and integrating over momentum space one obtains
@@
Thus, as in the non-relativistic case, this procedure leads to the same result as if one would directly substitute the Euler equations in order to write the time derivatives in terms of spatial gradients, and one obtains
@@
where $\sigma_{ab}$ is the symmetric traceless component, $\omega_{ab}$ the antisymmetric part and $u^{a}\frac{\partial\left(\quad\right)}{\partial x^{a}}:=\dot{\left(\quad\right)}$ denotes a proper time derivative, one is led to
@@ III.2 The projection method
As before one has $K^{\prime}_{R}=0$ if the Euler equations (5860) are imposed, which is justified since the expressions in parentheses in Eq. (72) are second order in this scheme (see Section IV). Therefore, eliminating $K^{\prime}_{R}$ when taking the orthogonal projection is not only consistent but also satisfies the necessary condition for the existence of the solution of Eq. (73), as explained in Section II.2. It is worthwhile commenting at this point that the relativistic calculation presents a major difference when compared with the non-relativistic counterpart. Not only do both procedures lead to different structures for the source term of the first-order Boltzmann equation, but the projection method does not eliminate the time derivatives. Consequently, they will be present in the solution for $\phi$, and thus they will also appear in the constitutive equations. From the point of view of classical irreversible thermodynamics this may be perceived as an inconsistency since dissipative fluxes are expected to be driven by spatial gradients of the state variables. However, in the relativistic framework, the constitutive equations resulting from Eq. (73) are not only valid but also equivalent to the ones obtained form Eq. (63) in Section III.1, in the sense of the general first-order theories. This is explained in detail in the next section.
@@ IV Discussion
In this work we have analyzed two different ways of implementing the CE method, which we referred to as the traditional procedure and the projection method. In the non-relativistic scenario both of them lead to the same integral equation, namely Eq. (22) and Eq. (26) together with Eq. (34). Meanwhile, in the relativistic case, the source terms differ from each other two versions of the integral equation for $\phi$ are obtained. Indeed, the traditional procedure leads to Eq. (63), whereas the projection method yields Eq. (73). In fact, both procedures constitute two different ways of imposing orthogonality of the source term of the integral equation to the kernel of the linearized collision operator, which is a necessary condition for the existence of the solution. This step is carried out by adjusting $\mathcal{D}$ in Section III.1 for the traditional method, and by projecting $p^{a}p^{b}$ onto the space orthogonal to $\ker(C)$ in Section III.2. Consequently, one would expect both equations to be valid and the resulting constitutive equations to be equivalent to each other within the linear scheme. Indeed, notice that Eq. (63) can be obtained from Eq. (73) and vice versa by imposing the Euler equations (59) and (60). The fact that one can go from one result to the other by using the Euler equations gives some insight into the kind of transformation that will ultimately relate the flux-force relations in both scenarios.
@@
$Q^{a}=\tilde{\kappa}_{1}\frac{{\rm D}^{a}n}{n}+\tilde{\kappa}_{2}\frac{{\rm D}^{a}T}{T}+\tilde{\kappa}_{4}\frac{q}{ch}F^{bc}u_{c},$ (74)
@@
    When coupled to the balance equations (42), do Eqs. (74-75) and/or Eqs. (76-77) lead to a physically sound theory for relativistic dissipative fluids regarding hyperbolicity, causality and stability?
The answers to these questions can be argued by considering the new proposals for first-order theories of relativistic dissipative fluids pK19 ; BDNK_PRX . In order to address the first three questions, in Appendix A we generalize the work by Kovtun pK19 to include the electromagnetic field and the additional freedom of adding terms proportional to the Euler equations (which are zero up to first order). This calculation shows that, even though $N^{a}$ and $T^{ab}$ can adopt the general form given in Eqs. (94-101), involving $18$ coefficients, only five linear combinations thereof are frame-invariant and independent of the representation, that is, independent of the addition of the Euler equations. Two of them are related to viscous dissipation and can be readily associated with the shear and bulk viscosities $\eta$ and $\zeta$, respectively, where $\eta$ is given in Eqs. (75,77) and
$\zeta=\pi_{1}+\frac{k_{B}}{c_{v}}\pi_{2}-\pi_{3},$ (78)
with $\pi_{i}$ the coefficients appearing in the general form for $P={\rm p}+\Pi$ in Eq. (98). The remaining three invariant quantities are related to the coefficients $\kappa_{i}$ appearing in the heat flux (see Eq. (100)) and are
@@
After these remarks, we are ready to provide answers to questions Q1–Q3. To this purpose, we first note that both sets of constitutive equations (Eqs. (74,75) and (76,77)) are particular cases of a general first-order theory in the Eckart frame, as described in Appendix A. Indeed, Eqs. (74,75) are obtained from Eqs. (98,100) with $\kappa_{i}=\tilde{\kappa}_{i}$ and $\pi_{i}=\tilde{\pi}_{i}$ where
@@
On the other hand, Eqs. (76,77) have $\kappa_{i}=\bar{\kappa}_{i}$ and $\pi_{i}=\bar{\pi}_{i}$ with
@@
Thus, regarding Q1 and Q2, one concludes that both representations are valid and equivalent in the linear case. Indeed, they are related to one another through Eqs. (107) and (108), which justifies the freedom of interchanging the acceleration term in Eq. (85) for a combination of the pressure gradient and the external electromagnetic force at any time using the Euler equation (Eq. (59)) to obtain Eq. (82) and vice versa. Moreover, it is easy to verify that condition (80) for the existence of global equilibria is satisfied in both cases.
@@
where $\vec{q}$ and $\overleftrightarrow{\pi}$ are the heat flux and viscous tensor respectively. [^1] Note that $\zeta=0$ for the Newtonian monoatomic ideal gas considered here. Whereas Eq. (88) has precisely the same form as its relativistic counterpart in Eq. (83), the relativistic generalization of the Fourier law is less straightforward. Indeed, if one were to inspect for the possibility of obtaining an expression similar to Eq. (87) in the relativistic scenario by setting $\kappa_{1}=\kappa_{3}=\kappa_{4}=0$ in Eq. (100) one would arrive at a contradiction since the compatibility condition (114) would imply that also $\kappa_{2}=0$. An alternative argument which allows one to establish the correct expression for the thermal conductivity is based on the structure of the entropy production, which is given by Eckart
@@
$\kappa=\frac{h^{2}}{\varepsilon}\mathcal{K}_{2}$ (90)
as the correct thermal conductivity in order to assure the non-equilibrium version of the second law is exactly satisfied in the absence of other dissipative fluxes. However, notice that the expressions for the force-flux couplings in Eqs. (82-83) and Eqs. (85-86) also lead to a positive entropy production when truncated to second order BDNK_PRX :
@@
where $\mathcal{E}$ is a correction term which, for the first set (Eqs. (82-83)) is
@@
and in the second case (Eqs. (85-86)) reads
@@
In both cases we see that, due to Eqs. (107,108), $\mathcal{E}$ is third order in the gradients, such that the entropy production is positive up to and including quadratic terms. [^2] The correction term $\mathcal{E}$ would be exactly zero for the ”mixed” option, Eqs. (85) and (83), corresponding to the Eckart theory; however this cannot be obtained within the CE approach depicted in this article.
Finally, to address Q4 one needs to resort to the stability and causality analysis of the system of equations (42) when coupled to the linear constitutive relations (82-83) or (85-86). Recent works point to the possibility of obtaining a hyperbolic and causal theory in a general frame and representation, provided a non-trivial set of inequalities for the coefficients appearing in the generalized constitutive equations is satisfied BDNK_PRX ; pK19 . The particular case of the theory obtained in Section III.1 has been previously analyzed and shown to lead to a stable system that does not violate causality in the comoving frame but becomes unstable in a general one reularubio . A thorough analysis of the theory involving constitutive equations (85-86) is beyond the scope of this article and will be addressed elsewhere.
Although for the sake of of clarity of the presentation we have restricted ourselves to the special-relativistic scenario in three space dimensions, our results can be easily generalized to a kinetic gas propagating on a curved spacetime of arbitrary dimensions.

### v2

@@ Abstract
Boltzmann equation, Chapman-Enskog approximation, relativistic fluids.
@@ I Introduction and preliminaries
In the relativistic scenario, the CE approximation was applied by pioneers in the field such as W. Israel Israel63 and S. R. de Groot deGrootLibro . However, the fact that its traditional form predicts first-order couplings between dissipative fluxes and spatial gradients of the state variables leads to its partial dismissal, based on the findings that such relations give rise to pathological theories for which the equilibrium configurations are generically unstable Hiscock . Recently, a new family of first-order theories has been proposed and shown to contain promising candidates to describe high temperature gases in curved spacetimes. These new theories allow one to consider a general frame and representation in which constitutive equations feature all possible couplings, including relations which involve time derivatives of the state variables. Under suitable restrictions, these theories have been shown to lead to physically sound equations BDNK_PRX ; pK19 . However, despite recent attempts bdnk19 ; rocha22 , their microscopical foundations are still not well understood from the point of view of the CE expansion. For recent progress based on the Hilbert expansion, see Ref. rHpK22 . The purpose of this work is to shed new light on this problem.
@@
The second, more formal, method follows the lines of Ref. LaurePaper (see also LaureLectureNotes ) by projecting the source term $L_{F}\left[\ln f^{\left(0\right)}\right]$ onto the orthogonal complement of the kernel of $C$, which consists of the collision invariants. Since the operator $C$ is self-adjoint when defined in an appropriate Hilbert space, this assures that the source term lies in the image of $C$.
@@ II Non-relativistic case
In order to compare the order of magnitude of each term in Boltzmann’s equation (1) one introduces dimensionless variables, denoted by overbars, as follows: $t=\left(L/v_{th}\right)\bar{t}$, $\vec{r}=L\vec{\bar{r}},\,$$\vec{v}=v_{th}\vec{\bar{v}}$, $g=\sqrt{2}v_{th}\bar{g}$, $f=\bar{f}/\left(v_{th}^{3}L^{3}\right)$, $\sigma=\pi d^{2}\bar{\sigma}$ and $\vec{F}=\left(mv_{th}^{2}/L\right)\vec{\bar{F}}$. Here $L$ is a characteristic macroscopic length scale, $v_{th}$ a typical mesoscopic speed and $d$ a microscopic scale of the order of magnitude of the molecules’ diameter. Defining the mean free path as $\ell:=L^{3}/\left(\sqrt{2}\pi d^{2}\right)$ and introducing these scaled variables in Eq. (1) one obtains (omitting the bars)
@@ II.1 The traditional Chapman-Enskog procedure
where the coefficients $\left(\mathcal{D}f\right)_{i}$ are still to be determined. This ansatz can be motivated by various arguments. The simplest one consists in noticing that the operator $\mathcal{D}$ involves only macroscopic variables, unlike the other two terms which depend on the molecular velocity. For a thorough discussion of this point the reader is referred to Refs. Ch-E and cercigniani1 for example.
@@
In order to determine the coefficients $\left(\mathcal{D}n\right)_{0}$, $\left(\mathcal{D}\vec{u}\right)_{0}$ and $\left(\mathcal{D}T\right)_{0}$, we multiply both sides of Eq. (17) by the vector of collision invariants $\vec{\psi}=\left(1,\vec{v},\frac{1}{2}mv^{2}\right)^{T}$ and $f^{\left(0\right)}$ and integrate over momentum space, obtaining
@@ III Relativistic Case
The order of magnitude for the terms in Eq. (1) can be assessed in a similar way as in the non-relativistic scenario with the following additional assumptions: $p^{a}=mv_{th}\bar{p}^{a}$, $f=\frac{1}{m^{3}v_{th}^{3}L^{3}}\bar{f}$, $c=v_{th}\bar{c}$, $q=e\bar{q}$, $F^{ab}=\frac{mv_{th}^{2}}{Le}\bar{F}^{ab}$, $\mathcal{F}=\sqrt{2}m^{2}v_{th}^{2}\bar{\mathcal{F}}$, where $e$ is the elementary charge. Proceeding as in Section II, one is led to the following dimensionless equation
@@
where we have once again omitted the bars over the dimensionless quantities in order to simplify the notation. Thus, in this scenario the collisional term also scales one order higher in $1/\text{Kn}$ than the transport part, which suggests the same expansion as in the previous case: $f=\sum_{i=0}^{\infty}\text{Kn}^{i}f^{\left(i\right)}$. As before, $f^{\left(0\right)}$ corresponds to a local equilibrium distribution function, parametrized by the dynamical variables $n$, $T$ and $u^{a}$ (where $u^{a}$ is the hydrodynamic four-velocity), and is given by the solution to $Q\left(f^{\left(0\right)}f^{\left(0\right)^{\prime}}\right)=0$ in Eq. (37), which leads to the Maxwell-Jüttner distribution function fJ11a ; CercignaniKremer-Book :
@@
However, out of equilibrium, there is no unique way to associate the components of $N^{a}$ and $T^{ab}$ to the state variables and fluxes in the system. In fact, there is a large family of valid representations as explained in Refs. pK19 ; BDNK_PRX . In the present work the so-called Eckart frame is considered. This particular choice is such that
@@ III.1 The relativistic generalization of the traditional Chapman-Enskog procedure
Notice that $u^{a}\frac{\partial f}{\partial x^{a}}$ in the first term on the right-hand side of Eq. (50) is the equivalent to the convective time derivative in the non-relativistic case. Thus it is reasonable to denote this term as $\mathcal{D}f$ and to write the relativistic Liouville operator as
@@
In order to find expressions for the coefficients $\left(\mathcal{D}n\right)_{0}$, $\left(\mathcal{D}T\right)_{0}$ and $\left(\mathcal{D}u^{c}\right)_{0}$ one can once again use the collision invariants. Indeed, multiplying Eq. (55) by $\vec{\psi}:=\left(1,p^{a}\right)^{T}$ and integrating over momentum space one obtains
@@
where $\dot{A}:=u^{a}\frac{\partial A}{\partial x^{a}}$ denotes the proper time derivative of a macroscopic quantity $A$, as measured by observers comoving with $u^{a}$. Thus, as in the non-relativistic case, this procedure leads to the same result as if one would directly substitute the Euler equations in order to write the time derivatives in terms of spatial gradients, and one obtains
@@
where $\sigma_{ab}$ is the symmetric traceless component and $\omega_{ab}$ the antisymmetric part, one is led to
@@ III.2 The projection method
As before one has $K^{\prime}_{R}=0$ if the Euler equations (5759) are imposed, which is justified since the expressions in parentheses in Eq. (71) are second order in this scheme (see Section IV). Therefore, eliminating $K^{\prime}_{R}$ when taking the orthogonal projection is not only consistent but also satisfies the necessary condition for the existence of the solution of Eq. (72), as explained in Section II.2. It is worthwhile commenting at this point that the relativistic calculation presents a major difference when compared with the non-relativistic counterpart. Both procedures lead to different structures for the source term of the first-order Boltzmann equation: in the traditional approach only spatial gradients and the electromagnetic force term $\frac{q}{c}F^{ba}u_{a}$ appear, whereas in the projection method the force term is absent. Moreover, in addition to the spatial derivatives, time derivatives of $T$ and $u^{a}$ survive the projection. Consequently, these time derivatives will be present in the solution for $\phi$, and thus they will also appear in the constitutive equations. From the point of view of classical irreversible thermodynamics this may be perceived as an inconsistency since dissipative fluxes are expected to be driven by spatial gradients of the state variables. However, in the relativistic framework, the constitutive equations resulting from Eq. (72) are not only valid but also equivalent to the ones obtained form Eq. (62) in Section III.1, in the sense of the general first-order theories. This is explained in detail in the next section.
@@ IV Discussion
In this work we have analyzed two different ways of implementing the CE method, which we referred to as the traditional procedure and the projection method. Both procedures constitute different ways of imposing orthogonality of the source term of the integral equation to the kernel of the linearized collision operator, which is a necessary condition for the existence of the solution. This step is carried out by adjusting $\mathcal{D}$ in Sections II.1 and  III.1 for the traditional method, and by explicitly projecting the source term onto the space orthogonal to $\ker(C)$ in Sections II.2 and  III.2. In the non-relativistic scenario both lead to the same integral equation, namely Eq. (22) and Eq. (26) together with Eq. (34). Meanwhile, in the relativistic case, the source terms differ from each other and two versions of the integral equation for $\phi$ are obtained: the traditional procedure leads to Eq. (62), whereas the projection method yields Eq. (72). One would expect both equations to be valid and the resulting constitutive equations to be equivalent to each other within the linear scheme. Indeed, notice that Eq. (62) can be obtained from Eq. (72) and vice versa by imposing the Euler equations (58) and (59). The fact that one can go from one result to the other by using the Euler equations gives some insight into the kind of transformation that will ultimately relate the flux-force relations in both scenarios.
@@
$Q^{a}=\tilde{\kappa}_{1}\frac{{\rm D}^{a}n}{n}+\tilde{\kappa}_{2}\frac{{\rm D}^{a}T}{T}+\tilde{\kappa}_{4}\frac{q}{ch}F^{ab}u_{b},$ (73)
@@
    When coupled to the balance equations (42), do Eqs. (73-74) and/or Eqs. (75-76) lead to a physically sound theory for relativistic dissipative fluids regarding hyperbolicity, causality and stability?
The answers to these questions can be argued by considering the new proposals for first-order theories of relativistic dissipative fluids pK19 ; BDNK_PRX . In order to address the first three questions, in Appendix A we generalize the work by Kovtun pK19 to include the electromagnetic field and the additional freedom of adding terms proportional to the Euler equations (which are zero up to first order). This calculation shows that, even though $N^{a}$ and $T^{ab}$ can adopt the general form given in Eqs. (95-102), involving $18$ coefficients, only five linear combinations thereof are frame-invariant and independent of the representation, that is, independent of the addition of the Euler equations. In order to describe these results, for definiteness we adopt the Eckart frame, in which case the first-order constitutive equations have the general form
$Q^{a}$ $=\kappa_{1}\frac{\mathrm{D}^{a}n}{n}+\kappa_{2}\frac{\mathrm{D}^{a}T}{T}+\kappa_{3}\frac{\dot{u}^{a}}{c^{2}}+\kappa_{4}\frac{q}{hc}F^{ab}u_{b},$ (77)
$\Pi$ $=\pi_{1}\frac{\dot{n}}{n}+\pi_{2}\frac{\dot{T}}{T}+\pi_{3}\theta,\qquad\tau^{ab}=-2\eta\sigma^{ab},$ (78)
with coefficients $\pi_{i}$ and $\kappa_{i}$. Out of the five frame- and representation-invariant quantities, two of them are related to viscous dissipation and can be readily associated with the shear and bulk viscosities $\eta$ and $\zeta$, respectively, where
$\zeta=\pi_{1}+\frac{k_{B}}{c_{v}}\pi_{2}-\pi_{3}.$ (79)
The remaining three invariant quantities are related to the coefficients $\kappa_{i}$ appearing in the heat flux and are
@@
After these remarks, we are ready to provide answers to questions Q1–Q3. To this purpose, we first note that both sets of constitutive equations (Eqs. (73,74) and (75,76)) are particular cases of Eqs. (77,78). Indeed, Eqs. (73,74) correspond to these equations with $\kappa_{i}=\tilde{\kappa}_{i}$ and $\pi_{i}=\tilde{\pi}_{i}$ where
@@
On the other hand, Eqs. (75,76) have $\kappa_{i}=\bar{\kappa}_{i}$ and $\pi_{i}=\bar{\pi}_{i}$ with
@@
Thus, regarding Q1 and Q2, one concludes that both representations are valid and equivalent in the linear case. Indeed, they are related to one another through the balance equations (see Eqs. (108) and (109)), which justifies the freedom of interchanging the acceleration term in Eq. (86) for a combination of the pressure gradient and the external electromagnetic force at any time using the Euler equation (Eq. (58)) to obtain Eq. (83) and vice versa. Moreover, it is easy to verify that condition (81) for the existence of global equilibria is satisfied in both cases.
@@
where $\vec{q}$ and $\overleftrightarrow{\pi}$ are the heat flux and viscous tensor respectively. [^1] Note that $\zeta=0$ for the Newtonian monoatomic ideal gas considered here. Whereas Eq. (89) has precisely the same form as its relativistic counterpart in Eq. (84), the relativistic generalization of the Fourier law is less straightforward. Indeed, if one were to inspect for the possibility of obtaining an expression similar to Eq. (88) in the relativistic scenario one would need to set $\kappa_{1}=\kappa_{3}=\kappa_{4}=0$ in Eq. (77). However, this would lead to a contradiction, since compatibility with global equilibrium would imply that also $\kappa_{2}=0$ (see Eq. (81)). An alternative argument which allows one to establish the correct expression for the thermal conductivity is based on the structure of the entropy production, which is given by Eckart
@@
$\kappa=\frac{h^{2}}{\varepsilon}\mathcal{K}_{2},$ (91)
as the correct thermal conductivity in order to assure the non-equilibrium version of the second law is exactly satisfied in the absence of other dissipative fluxes. However, notice that the expressions for the force-flux couplings in Eqs. (83-84) and Eqs. (86-87) also lead to a positive entropy production when truncated to second order BDNK_PRX :
@@
where $\mathcal{E}$ is a correction term which, for the first set (Eqs. (83-84)) is
@@
and in the second case (Eqs. (86-87)) reads
@@
In both cases we see that, due to balance equations (108,109), $\mathcal{E}$ is third order in the gradients, such that the entropy production is positive up to and including quadratic terms.222 The correction term $\mathcal{E}$ would be exactly zero for the ”mixed” option, Eqs. (86) and (84), corresponding to the Eckart theory; however this cannot be obtained within the CE approach depicted in this article.
Finally, to address Q4 one needs to resort to the stability and causality analysis of the system of equations (42) when coupled to the linear constitutive relations (83-84) or (86-87). Recent works point to the possibility of obtaining a hyperbolic and causal theory in a general frame and representation, provided a non-trivial set of inequalities for the coefficients appearing in the generalized constitutive equations is satisfied BDNK_PRX ; pK19 . The particular case of the theory obtained in Section III.1 has been previously analyzed and shown to lead to a stable system that does not violate causality in the comoving frame but becomes unstable in a general one reularubio . A thorough analysis of the theory involving constitutive equations (86-87) is beyond the scope of this article and will be addressed elsewhere.
Although for the sake of of clarity of the presentation we have restricted ourselves to the special-relativistic scenario in three space dimensions, our results can be easily generalized to a kinetic gas propagating on a curved spacetime of arbitrary dimensions. Notice that the results in Sections III.1 and III.2, regarding the different structure in the integral equations (62) and Eq. (72), are independent of the Eckart frame described at the beginning of that section. Therefore, it would be interesting to analyze whether our findings about the resulting differences in the constitutive equations (83-84) and (86-87), hold in frames more general than Eckart’s, based on the study of the homogeneous solution of the first-order equation similar to what has been performed in Ref. rHpK22 in the context of the Hilbert expansion.